### v1

# Dataset of GenAI-Assisted Information Problem Solving in Education

Abstract
Information Problem Solving (IPS) is a critical competency for academic and professional success in education, work, and life. The advent of Generative Artificial Intelligence (GenAI), particularly tools like ChatGPT, has introduced new possibilities for supporting students in complex IPS tasks. However, empirical insights into how students engage with GenAI during IPS and how these tools can be effectively leveraged for learning remain limited. Moreover, differences in background—shaped by cultural and socioeconomic factors—pose additional challenges to the equitable integration of GenAI in educational contexts. To address this gap, we present an open-source dataset collected from 279 students at a public Australian university. The dataset was generated through students’ use of FLoRA, a GenAI-powered educational platform that widely adopted in the field of learning analytics. Within FLoRA, students interacted with an embedded GenAI chatbot to gather information and synthesize it into data science project proposals. The dataset captures fine-grained, multi-dimensional records of GenAI-assisted IPS processes, including: (i) student–GenAI dialogue transcripts; (ii) writing process log traces; (iii) final project proposals with human-assigned assessment scores; (iv) surveys of biographic and prior knowledge in data science and AI; and (v) surveys capturing students’ GenAI experience and perceptions of GenAI’s effectiveness in supporting IPS. This dataset provides a valuable resource for advancing our understanding of GenAI’s role in educational IPS and informing the design of adaptive, inclusive AI-powered learning tools.
Background & Summary
Information Problem Solving (IPS) is a foundational competence in education, work, and everyday life, as people are increasingly required to identify information needs, locate relevant sources, critically evaluate information quality, and synthesise evidence across multiple resources to accomplish academic and professional tasks [1, 2]. Prior research has consistently shown that learners with stronger IPS skills achieve significantly better learning outcomes, including higher course grades [3, 4], improved knowledge retention, and greater effectiveness in subsequent workplace learning and performance [5]. Given the central role of IPS in academic success and lifelong learning, fostering learners’ capacity to engage in effective IPS has long been a core objective of educational research and practice.
To support empirical investigation and instructional design, IPS has been extensively operationalised as a multi-phase, goal-directed process that unfolds over time. Influential theoretical models conceptualise IPS as comprising a sequence of interrelated activities, including problem definition, information searching, information evaluation, information processing, and synthesis or presentation [6, 7, 8]. From a cognitive and self-regulated learning perspective, IPS is further understood as a dynamic process in which learners continuously monitor and regulate their strategies, decisions, and understanding while interacting with information sources and tools [9, 2].
In recent years, GenAI systems have been rapidly adopted by learners to support a wide range of educational activities, including IPS tasks [10]. Powered by large language models, GenAI tools enable learners to interact with information in conversational and iterative ways, thereby extending traditional forms of tool use in IPS. Rather than being confined to a single stage, GenAI can be integrated across multiple phases of the IPS process [11, 12]. During the early stages of IPS, learners may use GenAI to clarify task requirements, refine research questions, or decompose complex information problems into manageable sub-tasks. Beyond problem formulation, GenAI increasingly serves as an interactive channel for information gathering, allowing learners to obtain explanations, examples, and summaries through iterative dialogue rather than through conventional keyword-based search alone. In later phases, learners may rely on GenAI to summarise retrieved materials, compare perspectives across sources, generate integrative explanations, and obtain feedback on intermediate or final task products (e.g., essay drafts) to support evaluation and revision. Collectively, these practices indicate that GenAI functions as a flexible cognitive tool embedded throughout learners’ IPS workflows, with implications for both the nature and sequencing of IPS activities.
The rapid integration of GenAI into students’ learning practices has raised important questions about its potential impact on learning processes and outcomes. While GenAI offers new opportunities to support information access, sense-making, and feedback [13], concerns have also been expressed regarding over-reliance, reduced cognitive and metacognitive engagement, and exposure to inaccuracies and biases that may compromise information quality [14, 15, 16, 17]. These competing possibilities underscore the need for systematic empirical investigation into how learners actually make use of GenAI in their everyday IPS tasks. Although a growing body of research has begun to examine students’ use of GenAI in educational contexts, existing studies have often focused on isolated IPS phases, relied on controlled laboratory settings, or drawn primarily on learners’ retrospective self-reports, thereby providing only partial and indirect insights into real-world IPS practices. To address these limitations, we argue that a critical next step is to collect and openly share large-scale datasets that capture learners’ fine-grained behavioural traces as they engage in IPS processes with GenAI in authentic educational scenarios. Such datasets are essential for advancing data-driven understanding of GenAI-assisted IPS and for informing evidence-based instructional design and educational policy.
To address these challenges, this paper presents a large-scale, openly available dataset capturing learners’ IPS activities in an authentic GenAI-assisted educational context. The dataset was collected through a field study involving 519 postgraduate students at a public Australian university, who completed a two-week course assignment to develop a data science project proposal. Students were supported by FLoRA, a GenAI-powered system that integrates a GPT-4o chatbot with note-taking and writing tools and is designed to provide scaffolded support rather than direct solutions. The dataset comprises multiple complementary data modalities, including student–GenAI dialogue logs, fine-grained interaction traces of the writing process (e.g., mouse movements, keystrokes, and clicks), the project proposals produced as task outcomes together with instructor-assigned evaluation scores, and survey data capturing learners’ background characteristics and perceptions of GenAI use. By capturing learners’ interactions with GenAI as they gather information and construct project proposals in an authentic course setting, this dataset provides a unique empirical foundation for investigating IPS processes across phases and for developing data-driven instructional and AI-based scaffolding approaches to support effective student learning.
Methods
This section outlines the key stages involved in the preparation of the dataset, including ethical approval, data collection, anonymization, cleaning, qualitative coding of dialogues, and dataset release.
Ethical Approval and Informed Consent
This research project rigorously adheres to the ethical standards established by Monash University. The study was approved by Monash Human Research Ethics Committee under Project ID 42510. The dataset was collected from students in developing a data science project proposal as an assignment in an introductory level data science course for postgraduate students. Before commencing data collection, we distributed explanatory statements and forum announcements to all students to introduce the purpose of the study and to clarify that the data collected would be used exclusively for research purposes. Given that the study requires the use of FLoRA [18], we informed students that participation was optional and that they could complete the assignment without utilizing FLoRA if they preferred. We emphasised that all students had the right to withdraw from the research study at any time. Importantly, we assured them that their decision to participate or not would have no influence on their final assignment scores. By providing this information, we ensured that students could make informed decisions about their involvement, thereby upholding ethical research practices and respecting their autonomy.
Participants
The participants in this study comprised 519 postgraduate students and 65 students choose not to use FLoRA. These students had enrolled in the Foundations of Data Science course, either as an elective or a core component of their respective degrees. As this research was conducted as a field study, all participants worked within the same learning environment, ensuring consistency in the educational context.
Data Collection Procedure
Data collection was conducted using the FLoRA platform [18], which offers a suite of tools tailored for educational research, including an annotation tool, an AI-powered chatbot tool, and a writing tool. The study was systematically divided into four key phases: Pre-task survey, Training task, Main task, and Post-task survey. Each phase was designed to collect specific types of data. By capturing fine-grained, time-stamped logs of students’ annotations, trace of using the platform and interactions with the chatbot in real time, FLoRA yields high-resolution, process-oriented data on learners’ information problem-solving activities.
  1. Phase 1:
    Pre-task Survey. As noted above, students’ prior knowledge in the subject area and their previous experience with GenAI tools may significantly influence how they engage with GenAI during IPS tasks. To account for these factors, Phase 1 of the study involved administering two surveys. The first survey collected demographic information, such as students’ first-language background and academic major, to offer insights into their linguistic and educational contexts. The second survey focused on students’ prior knowledge and experience in data science and AI, including their familiarity with and usage of tools like ChatGPT. By collecting these background data, we aim to facilitate a deeper understanding of how individual differences may shape students’ interactions with GenAI in the context of IPS.
  2. Phase 2:
    Training. This phase was intentionally designed as a training stage to help students familiarise themselves with the FLoRA platform before engaging in the main task (i.e., Phase 3, described below). During this phase, students had full access to the same functionalities available in Phase 3, including the note-taking tool and the GenAI-powered chatbot. To facilitate effective onboarding, we provided comprehensive guidance along with illustrative examples demonstrating how to use each key feature of the platform. This preparatory step aimed to ensure that students became comfortable navigating FLoRA’s interface and functionalities, thereby enabling a smoother and more effective engagement with the main task in the subsequent phase.
  3. Phase 3:
    GenAI-Assisted Project Proposal Task. In Phase 3, students were required to develop a data science project proposal using the FLoRA platform, supported by an embedded GenAI-powered chatbot. They were first asked to select one topic from a set of 20 options spanning domains such as agriculture, education, finance, the game industry, healthcare, social media, and sports. Based on the chosen topic, students then had to formulate an idea for how data science could be applied in that domain and elaborate it into a written project proposal. Throughout this process, the chatbot functioned as an interactive research assistant, enabling students to ask questions, seek clarification, and obtain relevant information to support their information problem-solving activities. For example, a student exploring a topic on urban mobility might ask, “What are common data sources used in transportation studies?” or “Can you explain the concept of traffic flow prediction using machine learning?”. Also, students might use the chatbot to refine their ideas, test the relevance of their project scope, or request feedback on parts of their written proposal, such as asking, “Does this hypothesis seem focused enough?” or “Can you suggest improvements to this paragraph?” While the chatbot offered rich, responsive support, it was deliberately constrained from generating complete proposal content to ensure that students engaged meaningfully with the task and developed their own academic output. In this phase, trace data including the annotations students made, the interactions with the platform and the writing process logs was also recorded.
  4. Phase 4:
    Post-task Survey. Following the completion of the main task, participants were invited to complete a post-task survey aimed at evaluating their experiences with the FLoRA platform and the integrated GenAI support. This survey captured students’ perceptions of the usefulness and limitations of the chatbot, as well as their overall satisfaction with the learning experience. In particular, participants reflected on how the GenAI support influenced their ability to gather information and develop their project proposals. The feedback collected in this phase aim to provide insights into students’ engagement with GenAI, helping to inform the design of more effective and student-centred GenAI-assisted learning tools.
Data Preprocessing
To ensure participant privacy and comply with ethical standards, all data collected through the FLoRA platform underwent a thorough anonymisation process. Student identifiers, such as university email addresses used for logging into the platform, were replaced with randomly generated alphanumeric IDs. Additionally, all dialogues between students and the GenAI chatbot were manually reviewed to identify and remove any personally identifiable information, such as names, contact details, or references to specific individuals or locations. This manual review was critical for preserving participant confidentiality while maintaining the integrity and usefulness of the dialogue content for subsequent analysis.
The initial dataset comprised interaction records from 454 (total 519 and 65 did not use FLoRA) students who participated in the field study. To ensure analytical validity and data integrity, we applied a series of exclusion criteria to remove incomplete, inconsistent, or low-engagement cases. The following filters were applied:
  • Missing final proposal submission: Students who did not submit a final written project proposal via the FLoRA platform were excluded, as their task performance could not be evaluated.
  • Missing writing trace data: Students who used external tools (e.g., Microsoft Word or Google Docs) to compose their proposals instead of FLoRA’s built-in writing tool were removed, since writing logs, which are crucial for tracing IPS-related behaviours, could not be captured.
  • Chatbot interaction issues: Students who had no dialogue with the GenAI chatbot, or who submitted only a single prompt, were excluded. Manual review revealed that these minimal interactions were typically irrelevant or off-task (e.g., “Tell me a joke”) and did not reflect meaningful engagement with the IPS task. All dialogues not conducted in English were excluded to maintain consistency in linguistic and qualitative analysis.
  • Low similarity between FLoRA and Moodle submissions: Students were required to copy their FLoRA-composed proposal into a Microsoft Word document and submit it via Moodle for grading. To ensure consistency between the analysed proposal in FLoRA and the final assessed work submitted on Moodle, we computed similarity scores between both versions. Proposals with a similarity score below 0.85 (this threshold of 0.85 was determined through manual review of the lowest 10% most dissimilar cases) were excluded due to substantial post-transfer modifications.
After applying these exclusion criteria, the final dataset retained records from 279 students, each with complete system interaction trace data, annotation data, GenAI dialogues, writing process logs, final proposals, and proposal assessment scores, yielding a rich, high-quality dataset for investigating GenAI-supported IPS processes.
Dialogue Annotation
To systematically analyse how students interacted with the GenAI-powered chatbot, we developed a codebook for annotating student utterances (i.e., student intents) in Table 1. This codebook was based on established taxonomies for Conversational Information Seeking, especially the schemes developed in the work of Chu et al.[19] and Ren et al. [20]. However, while these prior schemes were designed for structured, goal-oriented search tasks, our study involved open-ended educational dialogues where students used GenAI to explore, clarify, and synthesise information for complex project-based learning. Consequently, we adapted and extended the original categories to reflect the broader range of communicative functions observed in our data.
Table 1: The codebook for labelling student utterances that show their intents.
Codes Description Examples
Idea Exploration Knowledge
Requests for factual, objective information with
clear, definitive answers.
Student: "What is a prediction model?"
Student: "What does a data scientist do?"
Judgement
Queries that require critical evaluation or reasoning,
often with multiple valid perspectives.
Student: "How can data science be applied in finance?"
Student: "What approach is most effective for this problem?"
Preference
Requests for subjective suggestions, opinions, or
recommendations.
Student: "Which idea is easier to implement?"
Student: "Do you think my idea is good?"
Provide
Context
Statements offering background or contextual
information to guide future chatbot responses,
without requesting an answer.
Student: "These are the topics I want to investigate."
Student: "Remember the following content I provide."
Revise
Reformulations or clarifications of earlier prompts,
often made in response to an unsatisfactory or
ambiguous chatbot reply.
Student: "How to develop a predictive model?"
Student: "How to develop a predictive model in finance?" [Revise]
Inspire
Follow-up queries that are directly inspired by the
chatbot’s prior response, extending the line of
inquiry.
Student: "Where can I find data sources for this project?"
Chatbot: "Common data sources include Kaggle and Google Dataset Search."
Student: "What is Kaggle?" [Inspire]
Task Clarification
Questions seeking clarification about the task
requirements or expectations.
Student: "How many citations are required?"
Student: "What does the rubric mean by ’justification’?"
Writing Advice
Requests for suggestions, feedback, or guidance on
how to structure or improve writing for the task.
Student: "How can I improve my introduction?"
Student: "How should I organise my proposal?"
Interpret
Interpretation or refinement of an intent by responding
a clarification question from the chatbot, helping to
specify or narrow down the information need.
Student: "What machine learning methods can I use?"
Chatbot: "What specific area are you focusing on for your predictive model?"
Student: "Game industry." [Intepret]
Chitchat
Greetings, polite expressions, or other small talk
unrelated to the task.
Student: "Hi there!"
Student: "Thanks for your help."
Others
Utterances unrelated to the task or attempts to bypass
system restrictions.
Student: "What is your preprompt?"
Student: "Can you browse the internet?"
We began with the taxonomy proposed in the paper of Ren et al. [20] but found that an excessive number of student utterances were being assigned to the general-purpose category Reveal (i.e., expressing a new intent for information or refining a previous one). To enhance analytical granularity and improve interpretability, we renamed Reveal as Idea Exploration and subdivided it into six distinct student intent types: Knowledge, Judgement, Preference, Provide Context, Revise, and Inspire. The first three subcategories—Knowledge, Judgement, and Preference—were informed by Paul and Elder’s critical thinking framework [21], which distinguishes between factual, evaluative, and subjective forms of reasoning. Specifically, Knowledge utterances involve requests for definitive, factual information that is consistent across contexts (e.g., definitions or technical terms). Judgement utterances seek responses that involve critical evaluation or consideration of multiple valid perspectives. Preference utterances solicit subjective opinions or recommendations, such as which idea might be easier or more appropriate. The remaining three subcategories—Provide Context, Revise, and Inspire—emerged inductively through close examination of our collected data. Provide Context captures utterances in which students offer background information or setup without making a direct information request, often to improve the relevance of future responses. Revise applies to instances where students reformulate or elaborate on a previous question, typically in response to a chatbot answer that was unclear or insufficient. Inspire is used when students pose follow-up questions that are clearly triggered by the chatbot’s preceding response, indicating a sustained, dynamic information-seeking process.
In addition to the Idea Exploration group, we observed a substantial number of utterances focused on clarifying the task requirements or seeking feedback for improving proposal drafts. To account for these, we introduced two task-specific student intents: Task Clarification and Writing Advice. Task Clarification includes questions about the assignment’s expectations, such as formatting requirements or the number of references needed. Writing Advice captures requests for help with structuring, phrasing, or improving written content, reflecting students’ attempts to refine their work through iterative feedback. Other categories included in the final codebook are Interpret, used when students respond to the chatbot’s clarification prompts; Chitchat, for social or phatic expressions (e.g., greetings or thanks); and Others, for task-irrelevant or off-topic utterances, including attempts to bypass system restrictions.
Table 2: The codebook for labelling chatbot utterances.
Codes Description Examples
Answer Fact
A direct response providing objective, factual information,
typically about a well-defined concept or term.
Student: "What is Infinity Ward’s net worth?"
Chatbot: "Infinity Ward’s net worth is 10 billion dollars."
Suggestion
A response offering examples, options, or guidance in a neutral
tone—without personal opinion or evaluative judgement.
Student: "Give me some interesting topics related to the Olympics dataset."
Chatbot: "Here are some interesting project topics: (a) Athlete Performance
Analytics of Medalists, (b) Athlete Injury Prediction and Prevention."
Critique
A subjective response expressing an evaluative opinion or
belief, without necessarily providing justification or evidence.
Student: "Do you think the topic of cardiovascular diseases is good?"
Chatbot: "The topic of cardiovascular diseases is interesting."
Clarify
A follow-up question or prompt from the chatbot requesting
additional information when the user’s intent is unclear or
underspecified.
Student: "Can you provide some references?"
Chatbot: "What areas of reference are you looking for?"
No Answer
A response indicating that no relevant information is available
or retrievable based on the user’s query.
Student: "Show me the product description for model FX-1."
Chatbot: "Sorry, no product has been found to satisfy your requirements."
Chitchat
Informal or social responses unrelated to task content, such as
greetings, acknowledgements, or conversational fillers.
Chatbot: "Sure, feel free to ask your new questions whenever you’re ready!"
Others
Responses unrelated to the IPS task or outside the intended
academic context.
Chatbot: "My GPT version is GPT-4o."
To analyse the communicative functions of the chatbot utterances, we also developed a response codebook grounded in the frameworks established in the paper of Chu et al. [19] and Ren et al. [20] for annotating the responses from conversational information systems. Building on these foundations, we refined and expanded the general Answer category into three specific subcodes: Fact, Suggestion, and Critique. These subcodes were designed to align conceptually with the three major cognitive categories used in the student intent codebook—namely, Knowledge, Judgement, and Preference. Specifically, Fact captures objective and verifiable information that remains consistent regardless of context. Suggestion refers to recommendations, examples, or procedural guidance presented without subjective bias. Critique, by contrast, reflects subjective judgments or evaluative opinions expressed without supporting justification. This distinction was necessary due to the generative nature of GPT-4o, which often produces multi-functional responses combining factual information, actionable suggestions, and personal-style commentary. To account for such complexity, chatbot utterances could be annotated with multiple labels when appropriate.
In addition to the refined answer types, we incorporated three codes from prior Conversational Information Seeking taxonomies [19, 20]: Clarify, No Answer, and Chitchat. Clarify denotes chatbot prompts that seek additional information or clarification of ambiguous student inputs. No Answer captures instances where the chatbot explicitly indicates that it cannot retrieve relevant information. Chitchat includes informal or social responses unrelated to the academic task. Finally, we introduced an Others category to classify responses that were irrelevant to the IPS task. These typically occurred when students posed off-topic or intentionally exploratory prompts, prompting the chatbot to reply beyond the intended task scope. Together, these two codebooks allow for a comprehensive and fine-grained analysis of student–GenAI interactions, enabling researchers to examine how students seek, process, and respond to information within authentic, GenAI-supported learning environments.
All dialogue data were annotated by three independent coders with relevant expertise in data science and educational technology. To ensure clarity and consistency in the annotation process, a detailed coding manual was developed, including operational definitions, decision rules, and illustrative examples for each code. This preparatory work was essential for aligning coders’ interpretations and reducing ambiguity during the annotation of both student intents and chatbot responses.
The annotation process began with an inter-rater calibration phase, during which all three coders independently annotated a randomly selected subset of 400 utterances—comprising both student and chatbot turns—from the full dataset. Three raters independently coded the data using a pairwise agreement method [22]. After completing the annotations, inter-rater agreement was evaluated using Cohen’s kappa. The average agreement score across all coder pairs was 0.92, indicating a high level of coding consistency. Discrepancies were then discussed collectively to refine the annotation guidelines and ensure shared understanding before proceeding with full dataset labelling. Following this calibration phase, the full dataset was divided into three subsets, with each coder responsible for annotating two of the three. This ensured that every utterance was independently coded by two coders. When coding disagreements arose, the third coder served as an adjudicator to resolve them through discussion and reach consensus. This dual-coding plus adjudication approach strengthened the reliability and validity of the final annotated dataset, supporting rigorous analysis of student–GenAI interactions in the context of IPS tasks.
In total, the annotated dataset comprises 3,560 student prompts and chatbot responses, each labelled according to the developed codebooks described in Table 1 and Table 2, with the distribution of intent and response types visualised in Figure 1 and the distribution student activity visualised in Figure 2.
Figure 1: Students Questions and Chatbot Answers Category Distribution
Figure 2: Students Activity Distribution
Data Records
The dataset is publicly available on GitHub [^1] https://github.com/Xinyu-Li/FLoRA_Open_Data_Monash_FIT5145 under the BSD-2-Clause license. It comprises multiple CSV-format data files, as detailed below.
Dialogue Data
This file contains all conversations between students and the GenAI-powered chatbot. Each row corresponds to a single exchange, comprising a student utterance and the corresponding chatbot response. It includes the following columns:
  • user_id — A pseudonymised unique identifier for each student.
  • user_ask_time — The timestamp when the student submitted the utterance to the chatbot.
  • user_utterance_text — The content of the student’s utterance.
  • user_utterance_code — The assigned code representing the student’s intent, labelled according to the codebook defined in Table 1.
  • chatbot_utterance_time — The timestamp when the chatbot provided its response.
  • chatbot_utterance_text — The textual content of the chatbot’s response. This field is stored in JSON format as a sequential list of text segments, with each segment may be associated with a distinct chatbot_utterance_code (described below).
  • chatbot_utterance_code — The corresponding list of codes (also in JSON format) representing the response types for each text segment in chatbot_utterance_text, based on the codebook defined in Table 2.
System Log Traces
This file contains the system-generated log traces recorded during students’ interactions with the FLoRA platform. Each row corresponds to a single log event and includes the following columns:
  • user_id — A pseudonymised identifier for the student who generated the log event.
  • log_trace_time - The timestamp indicating when the log event occurred.
  • url - An identifier for the webpage where the log event took place. This may include pages where students used FLoRA’s sidebar tools (e.g., the Annotation tool or the GenAI-powered chatbot).
  • learning_action_label - A high-level label describing the student’s learning action, automatically mapped from low-level log data. For example, keystrokes made while drafting a project proposal are labelled as WRITE_ESSAY. A complete list of action labels and their definitions is provided in Table 3.
  • action_category_label — A broader category that groups related learning_action_label entries. For instance, WRITE_ESSAY, PASTE_TEXT_ESSAY, SAVE_ESSAY, OPEN_ESSAY, and CLOSE_ESSAY are all classified under the Essay action category. Descriptions for all action categories can also be found in Table 3.
Table 3: The learning action categories
No. Action No. Sub_Action Sub Action Definition
0 START 0.1 START_TASK Indicate when learners start the task.
1 INSTRUCTION 1.2 TASK_REQUIREMENT Learners open task requirement page (task instruction page) to read about what the task requires them to do.
2 READING 2.1 READING_MATERIALS Learners re-read relevant reading materials.
2.2 REREADING_MATERIALS Learners re-read relevant reading materials.
3 ESSAY 3.1 OPEN_ESSAY Learners open the essay and read their essay without new writing.
3.2 WRITE_ESSAY Learners open the essay and write based on reading materials.
3.3 PASTE_TEXT_ESSAY Learners copy and paste materials from reading content to the essay window.
3.4 SAVE_ESSAY Learners click the save button to save the essay.
3.5 CLOSE_ESSAY Learners close the essay window.
4 ANNOTATION 4.1 CREATE_NOTE Learners create notes from reading materials.
4.2 CREATE_HIGHLIGHT Learners create highlights on the reading materials.
4.3 READ_ANNOTATION Learners click on annotations or open the annotation tool to read their notes or highlights.
4.4 LABEL_ANNOTATION Learners label or add new labels, or accept suggested labels on their notes or highlights.
4.5 EDIT_ANNOTATION Learners edit their annotations, such as edit one note on one keyword.
4.6 DELETE_ANNOTATION Learners delete their annotations, such as delete the highlight on one sentence.
4.7 SEARCH_ANNOTATION Learners use search annotation tool to search and check their annotations.
4.8 CLOSE_ANNOTATION Learners close the annotation window.
5 CHATBOT 5.1 OPEN_CHATBOT Learners open the chatbot window.
5.2 ASK_CHATBOT Learners type question and send questions to chatbot.
5.3 READ_FEEDBACK_CHATBOT Learners move mouse, scroll mouse or select text in the chatbot response content.
5.4 CLOSE_CHATBOT Learners close the chatbot window.
6 NAVIGATION 6.1 PAGE_NAVIGATION Learners navigate through several pages (stay less than 6 seconds)
6.2 TABLE_OF_CONTENT Learners check the table of content, such as scrolling in that area
6.3 TRY_OUT_TOOLS Learners quickly (less than 3 seconds) open and close tools for the first time without using them
7 OFF_TASK 7.1 OFF_TASK Learners being inactivity for more than 5 minutes
There are two types of web pages in FLoRA, namely instructions and inspirational_materials. The former refers to the only web page detailing the assignment instruction; the latter refers to various reading materials provided by the teachers for students to get inspiration about different topics.
Writing logs
The writing log data contains all the keystroke events and the writing content that happen while students writing the proposal. It contains the following columns:
  • user_id - the unique identifier for each student.
  • writing_log_time - the timestamp this writing log trace is recorded.
  • url - the web page URL this trace is recorded.
  • writing_content - the writing text that is recorded after each keystroke.
Annotation logs
The annotation log data contains all the annotation content that happen while students writing the proposal. It contains the following columns:
  • user_id - the unique identifier for each student.
  • annotation_log_time - the timestamp this annotation log trace is recorded.
  • url - the web page URL this trace is recorded.
  • highlight_text - the text content that is highlighted.
  • default_tag - the default labels given to the annotation when students label some text.
  • customised_tags - the extra tags students created by themselves.
  • notes_text - the notes students given to each annotation.
Proposal products and scores
The proposal data contains all students’ final proposals and scores. It contains the following columns:
  • user_id - the unique identifier for each student.
  • final_proposal - the proposal text students finished.
  • proposal_score - the proposal score given by the Data Science course tutors.
Pre-task biographic survey
This table contains the pre-task biographic survey questions and responses. It contains the following columns:
  • user_id - the unique identifier for each student.
  • response_time - time when student finish this survey.
  • gender - the gender of students.
  • age - the age of students.
  • mother_tongue - the mother tongue of students.
  • degree - whether students are bachelor or master degree.
  • degree_level - which year students are in their degree.
  • major - the major of students.
Pre-task prior knowledge and experience survey
This table contains the pre-task prior knowledge and experience survey questions and responses. It contains the following columns:
  • user_id - the unique identifier for each student.
  • response_time - time when student finish this survey.
  • has_prior_knowledge_in_DS - whether students has prior knowledge and experience in data science or a related field (e.g., data mining).
  • years_of_experience - how many years of experience do students have.
  • has_prior_knowledge_in_genai - whether have any prior knowledge and experience in generative AI tools such as ChatGPT and GPT-4.
  • extent_of_familiar_genai - to what extent students think they are familiar with generative AI tools.
Post-task platform usage surveys
This table contains the post-task platform usage survey questions and responses. It contains the following columns:
  • user_id - the unique identifier for each student.
  • response_time - time when student finish this survey.
  • extent_of_materials_in_flora_useful - to what extent students think the provided materials on FLoRA (e.g., those documents discussing the application of data science in different domains) are useful to find a project topic.
  • extent_chatbot_useful_for_finding_topic - to what extent students think the GPT-powered chatbot on FLoRA is useful to find a project topic.
  • extent_chatbot_useful_for_finish_assignment - to what extent students think the GPT-powered chatbot on FLoRA is useful to accomplish the assignment.
  • extent_engagement - to what extent students describe their engagement in using the GPT-powered chatbot to accomplish the assignment.
  • explanation - the reasons why students did not engage with the GPT-powered chatbot.
  • using_gpt_enhance_interest - whether students think interacting with the GPT-powered chatbot enhanced their interest in the subject matter.
  • extent_understanding_of_subject_improved - to what extent has students’ understanding of the subject matter improved after using the FLoRA platform.
  • satisfy_flora - whether students are satisfied with the experience on the FLoRA platform.
  • use_other_tools - the other tools students choose the finish the assignment.
Technical Validation
To ensure the reliability and validity of all data used in this study, we implemented a multi‑step technical validation protocol. Each step addresses a different potential threat to data quality, and together they provide a coherent basis for confidence in both the dataset and the findings derived from it.
First, we verified the data collection platform (instrumentation validity). FLoRA [18] is open-source and has been extensively validated in prior research. For example, FLoRA has been utilised across multiple peer-reviewed journal articles [23, 24, 25, 26], and conferences publications [27, 28]. Moreover, data gathered via FLoRA has been validated using think-aloud protocols and eye-tracking data, demonstrating robust methodological rigour and thought [29, 30]. On this basis, we adopted FLoRA as the data collection system for this study.
Figure 3: Students Interaction Trace Data Distribution
Second, we tested the stability of the trace data across cohorts (temporal stability and replicability). We compared a pilot dataset collected under identical experimental conditions in Semester One 2024 (2024S1) with the main dataset from Semester Two 2024 (2024S2), which we intend to release as open data. After conducting normality and homogeneity‑of‑variance tests, we found that both datasets were non‑normal but exhibited equal variances. We therefore used a Mann–Whitney U test, which yielded p = 0.171063 (> 0.05), indicating no significant difference in the distribution of trace logs per student between 2024S1 and 2024S2. Figure 3 shows comparable right‑skewed distributions across semesters, with similar medians and a comparable spread of outliers; although 2024S2 includes more students, the distributional patterns remain stable. This consistency supports the reliability of the collection process and the generalisability of our findings.
Third, we quality‑checked all utterance data (content integrity). Trained data raters examined every user question and chatbot response line by line to verify appropriateness (removing any offensive language), relevance to the study context and no non-English content. This review ensured that anomalies or inappropriate items were identified and addressed promptly.
Fourth, we validated the assessment data (assessment reliability). Expert teachers evaluated all submitted essays using detailed, criterion‑based rubrics. Their discipline expertise and consistent application of the criteria ensured rigorous grading and supported reliable links between learning outcomes and interactions with the GenAI system.
Finally, we verified the survey data (background variable integrity). The research team reviewed all pre‑task and post‑task responses, removing incomplete entries when appropriate. This step ensured that the background information and participant feedback used in subsequent analyses were reliable.
Together, these procedures triangulate platform validity, cross‑semester stability, dialogue quality, assessment reliability, and survey integrity, thereby strengthening the overall credibility of the dataset and the conclusions drawn from it.
Usage Notes
This dataset is best suited for research about IPS, as well as for detailed writing process analytics. Researchers can use the dialogues to infer planning, monitoring, evaluation, and adaptation processes across sessions. Because the interaction log data is pre‑labelled with SRL processes in FLoRA (using Bannert’s SRL framework [31, 29]), this data can also be used for SRL analysis to evaluate students cognition and metacognition activities during the learning.
Building on these process measures, the dataset also enables rich GenAI interaction studies and performance modelling. Prompting strategies, query refinement, and response uptake can be characterized using dialogue features such as turn length, intent categories, reformulation frequency, and evaluation moves, which can then be connected to subsequent writing actions in the trace data. These micro‑level behaviours, along with GenAI usage patterns, can be related to macro‑level results such as proposal scores, while controlling for prior knowledge and background survey variables.
Finally, the data support comparative evaluations of GenAI’s impact and survey‑based psychometric analyses, which complement the process‑level work. Because GenAI adoption is self‑selected, contrasts between users and non‑users, or between low‑ and high‑intensity users, are best examined with quasi‑experimental designs (e.g., propensity score matching, inverse probability weighting, difference‑in‑differences around key milestones). In addition, survey responses on GenAI effectiveness, IPS self‑efficacy, and related constructs can be analysed psychometrically, testing measurement invariance across demographic groups before relating survey factors to trace‑ and dialogue‑based behaviours.
Code Availability
The FLoRA that is utilised for data collection in this study was developed by our team and is available as open-source software on GitHub. A comprehensive README file is included to provide detailed instructions on deploying and using the platform for data collection. The platform supports a diverse range of data types and experimental scenarios. By adjusting configurations, users can tailor the platform to collect the specific data they require. The data cleaning code and the data are open sourced at another GitHub repository.
Acknowledgements
This research is funded partially by the Australian Government through the Australian Research Council (project number DP240100069 and DP220101209) and the Jacobs Foundation (CELLA 2 CERES).
Author contributions statement
X.L., K.Y., Y.C., L.L., and Z.L. selected the data and conducted data labelling. X.L., K.Y., J.W. and G.C. prepared the dataset for publication. G.C. and D.G. conceptualised and supervised the data collection. X.L. drafted the manuscript. K.Y., J.W., Y.C., G.C. and D.G. reviewed and approved the final version of the manuscript.
Competing interests
The authors declare no competing interests.

### v2

@@
 To address this gap, we present an open-source dataset collected from 279 students at a public Australian university.
-The dataset was generated through students’ use of FLoRA, a GenAI-powered educational platform that widely adopted in the field of learning analytics.
+The dataset was generated through students’ use of FLoRA, a GenAI-powered educational platform that is widely adopted in the field of learning analytics.
 Within FLoRA, students interacted with an embedded GenAI chatbot to gather information and synthesize it into data science project proposals.
-The dataset captures fine-grained, multi-dimensional records of GenAI-assisted IPS processes, including: (i) student–GenAI dialogue transcripts; (ii) writing process log traces; (iii) final project proposals with human-assigned assessment scores; (iv) surveys of biographic and prior knowledge in data science and AI; and (v) surveys capturing students’ GenAI experience and perceptions of GenAI’s effectiveness in supporting IPS.
+The dataset captures fine-grained, multi-dimensional records of GenAI-assisted IPS processes, including: (i) student–GenAI dialogue transcripts; (ii) writing process log traces; (iii) final project proposals with human-assigned assessment scores; (iv) two surveys assessing students demographic background and their prior knowledge and experience in data science and AI; and (v) surveys capturing students’ perceptions of GenAI’s effectiveness in supporting IPS and platform use experience.
 This dataset provides a valuable resource for advancing our understanding of GenAI’s role in educational IPS and informing the design of adaptive, inclusive AI-powered learning tools.
@@
 Rather than being confined to a single stage, GenAI can be integrated across multiple phases of the IPS process [44, 4].
-During the early stages of IPS, learners may use GenAI to clarify task requirements, refine research questions, or decompose complex information problems into manageable sub-tasks.
-Beyond problem formulation, GenAI increasingly serves as an interactive channel for information gathering, allowing learners to obtain explanations, examples, and summaries through iterative dialogue rather than through conventional keyword-based search alone.
-In later phases, learners may rely on GenAI to summarise retrieved materials, compare perspectives across sources, generate integrative explanations, and obtain feedback on intermediate or final task products (e.g., essay drafts) to support evaluation and revision.
+During the early stages of IPS, learners may use GenAI to clarify task requirements, refine research questions, or decompose complex information problems into manageable sub-tasks [22, 8].
+Beyond problem formulation, GenAI increasingly serves as an interactive channel for information gathering, allowing learners to obtain explanations, examples, and summaries through iterative dialogue rather than through conventional keyword-based search alone [14].
+In later phases, learners may rely on GenAI to summarise retrieved materials, compare perspectives across sources, and generate integrative explanations [16], and obtain feedback on intermediate or final task products (e.g., essay drafts) to support evaluation and revision [7, 43].
 Collectively, these practices indicate that GenAI functions as a flexible cognitive tool embedded throughout learners’ IPS workflows, with implications for both the nature and sequencing of IPS activities.
@@
 These competing possibilities underscore the need for systematic empirical investigation into how learners actually make use of GenAI in their everyday IPS tasks.
-Although a growing body of research has begun to examine students’ use of GenAI in educational contexts, existing studies have often focused on isolated IPS phases, relied on controlled laboratory settings, or drawn primarily on learners’ retrospective self-reports, thereby providing only partial and indirect insights into real-world IPS practices.
-To address these limitations, we argue that a critical next step is to collect and openly share large-scale datasets that capture learners’ fine-grained behavioural traces as they engage in IPS processes with GenAI in authentic educational scenarios.
+Although a growing body of research has begun to examine students’ use of GenAI in educational contexts, existing studies have often focused on isolated IPS phases [30, 23], relied on controlled laboratory settings [34], or drawn primarily on learners’ retrospective self-reports [28], thereby providing only partial and indirect insights into real-world IPS practices [36, 24].
+To address these limitations, a critical next step is to collect and openly share large-scale datasets that capture learners’ fine-grained behavioural traces as they engage in IPS processes with GenAI in authentic educational scenarios.
 Such datasets are essential for advancing data-driven understanding of GenAI-assisted IPS and for informing evidence-based instructional design and educational policy.
@@
 The dataset was collected through a field study involving 519 postgraduate students at a public Australian university, who completed a two-week course assignment to develop a data science project proposal.
-Students were supported by FLoRA, a GenAI-powered system that integrates a GPT-4o chatbot with note-taking and writing tools and is designed to provide scaffolded support rather than direct solutions.
+Students were supported by FLoRA, a GenAI-powered system that integrates a GPT-4o chatbot with note-taking and writing tools and is designed to provide scaffolded support rather than direct solutions [20, 19].
 The dataset comprises multiple complementary data modalities, including student–GenAI dialogue logs, fine-grained interaction traces of the writing process (e.g., mouse movements, keystrokes, and clicks), the project proposals produced as task outcomes together with instructor-assigned evaluation scores, and survey data capturing learners’ background characteristics and perceptions of GenAI use.
@@
 This research project rigorously adheres to the ethical standards established by Monash University.
-The study was approved by Monash Human Research Ethics Committee under Project ID 42510.
+The study was approved by the Monash University Human Research Ethics Committee under Project ID 42510.
 The dataset was collected from students in developing a data science project proposal as an assignment in an introductory level data science course for postgraduate students.
 Before commencing data collection, we distributed explanatory statements and forum announcements to all students to introduce the purpose of the study and to clarify that the data collected would be used exclusively for research purposes.
-Given that the study requires the use of FLoRA [18], we informed students that participation was optional and that they could complete the assignment without utilizing FLoRA if they preferred.
+Given that the study required the use of FLoRA [19], we informed students that participation was optional and that they could complete the assignment without utilizing FLoRA if they preferred.
 We emphasised that all students had the right to withdraw from the research study at any time.
@@
 As this research was conducted as a field study, all participants worked within the same learning environment, ensuring consistency in the educational context.
+Data were collected in two offerings of the same course: a pilot cohort in Semester 1, 2024 (2024S1) and the main cohort in Semester 2, 2024 (2024S2).
+Unless otherwise stated, the open dataset released with this paper corresponds to 2024S2; 2024S1 was used only for stability checks reported in the Technical Validation section.
 Data Collection Procedure
-Data collection was conducted using the FLoRA platform [18], which offers a suite of tools tailored for educational research, including an annotation tool, an AI-powered chatbot tool, and a writing tool.
+Data collection was conducted using the FLoRA platform [19], which offers a suite of tools tailored for educational research, including an annotation tool, an GenAI-powered chatbot tool, and a writing tool.
+These tools were selected to capture fine-grained information problem solving traces across key phases—students’ in-text highlighting/notes while engaging with abundant readings (annotation tool), their iterative querying and evaluation actions during AI-assisted search and reasoning (GenAI-powered chatbot), and their step-by-step proposal drafting and revision logs (writing tool).
+In particular, the built-in writing tool captured keystroke-level writing logs and intermediate draft states, enabling fine-grained analyses of drafting and revision as part of the synthesis/production phase of IPS.
 The study was systematically divided into four key phases: Pre-task survey, Training task, Main task, and Post-task survey.
 Each phase was designed to collect specific types of data.
-By capturing fine-grained, time-stamped logs of students’ annotations, trace of using the platform and interactions with the chatbot in real time, FLoRA yields high-resolution, process-oriented data on learners’ information problem-solving activities.
+FLoRA recorded fine-grained, time-stamped process data while students engaged in reading, annotating, chatting with the GenAI assistant, and writing.
+These data included (i) students’ annotations and notes, (ii) student–chatbot interaction logs, and (iii) time-stamped trace events generated during platform use, such as navigation events and clicks, mouse interactions (e.g., move/click/scroll), and keyboard interactions (keystrokes).
+Each raw event was stored as a log entry with timestamps and interface context (e.g., the active page/tool).
+Together, these time-stamped logs yield high-resolution, process-oriented data on learners’ information problem-solving activities.
   1. Phase 1:
@@
     GenAI-Assisted Project Proposal Task.
-In Phase 3, students were required to develop a data science project proposal using the FLoRA platform, supported by an embedded GenAI-powered chatbot.
+In Phase 3, students were required to develop a data science project proposal using the FLoRA platform, supported by an embedded GenAI-powered chatbot (GPT-4o).
 They were first asked to select one topic from a set of 20 options spanning domains such as agriculture, education, finance, the game industry, healthcare, social media, and sports.
-Based on the chosen topic, students then had to formulate an idea for how data science could be applied in that domain and elaborate it into a written project proposal.
+Based on the chosen topic, students then were asked to formulate an idea for how data science could be applied in that domain and elaborate it into a written project proposal.
 Throughout this process, the chatbot functioned as an interactive research assistant, enabling students to ask questions, seek clarification, and obtain relevant information to support their information problem-solving activities.
-For example, a student exploring a topic on urban mobility might ask, “What are common data sources used in transportation studies?” or “Can you explain the concept of traffic flow prediction using machine learning?”.
-Also, students might use the chatbot to refine their ideas, test the relevance of their project scope, or request feedback on parts of their written proposal, such as asking, “Does this hypothesis seem focused enough?” or “Can you suggest improvements to this paragraph?” While the chatbot offered rich, responsive support, it was deliberately constrained from generating complete proposal content to ensure that students engaged meaningfully with the task and developed their own academic output.
+For example, a student exploring a topic on urban mobility could ask, “What are common data sources used in transportation studies?” or “Can you explain the concept of traffic flow prediction using machine learning?”.
+Also, students could use the chatbot to refine their ideas, test the relevance of their project scope, or request feedback on parts of their written proposal, such as asking, “Does this hypothesis seem focused enough?” or “Can you suggest improvements to this paragraph?” While the chatbot offered rich, responsive support, it was deliberately constrained from generating complete proposal content to ensure that students engaged meaningfully with the task and developed their own academic output.
 In this phase, trace data including the annotations students made, the interactions with the platform and the writing process logs was also recorded.
@@
 The feedback collected in this phase aim to provide insights into students’ engagement with GenAI, helping to inform the design of more effective and student-centred GenAI-assisted learning tools.
+In addition to these process logs, we collected students’ final proposals, rubric-based assessment scores, and survey responses as outcome and background measures.
 Data Preprocessing
@@
   • Missing final proposal submission: Students who did not submit a final written project proposal via the FLoRA platform were excluded, as their task performance could not be evaluated.
-  • Missing writing trace data: Students who used external tools (e.g., Microsoft Word or Google Docs) to compose their proposals instead of FLoRA’s built-in writing tool were removed, since writing logs, which are crucial for tracing IPS-related behaviours, could not be captured.
+  • Missing writing trace data: Students who used external tools (e.g., Microsoft Word or Google Docs) to compose their proposals instead of FLoRA’s built-in writing tool were removed, since writing logs including keystrokes, which are crucial for tracing IPS-related behaviours, could not be captured.
   • Chatbot interaction issues: Students who had no dialogue with the GenAI chatbot, or who submitted only a single prompt, were excluded.
@@
 Proposals with a similarity score below 0.85 (this threshold of 0.85 was determined through manual review of the lowest 10% most dissimilar cases) were excluded due to substantial post-transfer modifications.
+For similarity score calculation, we utilised the cosine similarity method.
+Specifically, we first converted the textual data into high-dimensional dense vectors using OpenAI’s text-embedding-3-small model.
+This model maps text into a semantic vector space, capturing contextual meaning more effectively than traditional keyword-based methods (e.g., TF-IDF).
+The cosine similarity score was then computed between the resulting embedding vectors to determine the degree of semantic alignment.
 After applying these exclusion criteria, the final dataset retained records from 279 students, each with complete system interaction trace data, annotation data, GenAI dialogues, writing process logs, final proposals, and proposal assessment scores, yielding a rich, high-quality dataset for investigating GenAI-supported IPS processes.
+Students’ final project proposals were graded as part of the normal course assessment by the course teaching staff, all of whom had multiple years of teaching experience.
+Grading followed a criterion-based rubric.
+The rubric criteria included: Clear description of the goals of the project (3%), Appropriateness of topic (3%), Clear description of the business benefits and challenges (3%), Novelty/Creativity (3%), and Overall clarity of the report (3%).
+As this was a field study embedded in authentic teaching practice, grading was conducted following standard course procedures (i.e., no additional double-marking or sampling-based moderation was introduced solely for research purposes).
+Survey responses were linked to the main dataset using students’ anonymised IDs and filtered to retain only the final analytical sample (i.e., the same 279 students retained after the interaction-log and proposal-based exclusion criteria).
+Responses were screened and incomplete entries for key variables were removed when appropriate.
+The complete survey instruments (all questions and response options) are provided in Supplementary Information Table S1, S2, and S3.
+Learning Action Logs and Action Coding
+FLoRA automatically parsed raw trace events into interpretable learning-action labels (Table 1) using a rule-based mapping, enabling interpretable analyses of students’ IPS behaviours [12].
+Such trace data provide an unobtrusive record of learners’ moment-to-moment activities and are valuable for studying information problem solving (IPS), because IPS involves not only asking questions to the chatbot but also iterative reading, revisiting sources, writing, and tool-mediated exploration that unfold as observable interaction sequences.
+These traces are important for IPS research because they capture not only information seeking and evaluation behaviours, but also the synthesis/production activities (i.e., writing and revision) through which learners integrate information into an academic proposal.
+In the current dataset, raw trace events were automatically mapped into learning action categories using a rule-based trace parser (i.e., an action library approach).
+Specifically, raw events were mapped into interpretable learning action categories (Table 1) based on the interface context and a small set of deterministic rules: (i) page context rules (e.g., events on the task-instruction interface were labelled as TASK_REQUIREMENT); (ii) reading vs.
+re-reading rules based on repeated visits to the same reading page (first visit labelled as READING_MATERIALS, subsequent visits labelled as REREADING_MATERIALS); (iii) navigation filtering based on dwell time (rapid page switches with dwell time << 6 seconds were labelled as PAGE_NAVIGATION); and (iv) tool- and input-based rules for writing/annotation/chatbot actions (e.g., opening/closing tools, saving, keystroke-based writing, and paste events such as PASTE_TEXT_ESSAY).
+Periods of inactivity longer than 5 minutes were labelled as OFF_TASK.
+Importantly, we did not directly assign self-regulated learning (SRL) process labels (e.g., metacognitive planning/monitoring or cognitive reading/elaboration) in this study.
+Readers interested in SRL process-level coding can derive such labels by applying a process library that maps recurring sequences of learning actions to SRL processes aligned with Bannert’s model [1].
+We refer readers to Fan et al.
+[12] for the validated action/process library framework and examples of how action sequences are mapped to SRL processes.
+Table 1: The learning action categories
+No. Action No. Sub_Action Sub Action Definition
+0 START 0.1 START_TASK Indicate when learners start the task.
+1 INSTRUCTION 1.2 TASK_REQUIREMENT Learners open task requirement page (task instruction page) to read about what the task requires them to do.
+2 READING 2.1 READING_MATERIALS Learners read relevant reading materials (first visit within a session).
+2.2 REREADING_MATERIALS Learners re-read relevant reading materials (revisit of a previously visited page).
+3 ESSAY 3.1 OPEN_ESSAY Learners open the essay and read their essay without new writing.
+3.2 WRITE_ESSAY Learners open the essay and write based on reading materials.
+3.3 PASTE_TEXT_ESSAY Learners copy and paste materials from reading content to the essay window.
+3.4 SAVE_ESSAY Learners click the save button to save the essay.
+3.5 CLOSE_ESSAY Learners close the essay window.
+4 ANNOTATION 4.1 CREATE_NOTE Learners create notes from reading materials.
+4.2 CREATE_HIGHLIGHT Learners create highlights on the reading materials.
+4.3 READ_ANNOTATION Learners click on annotations or open the annotation tool to read their notes or highlights.
+4.4 LABEL_ANNOTATION Learners label or add new labels, or accept suggested labels on their notes or highlights.
+4.5 EDIT_ANNOTATION Learners edit their annotations, such as edit one note on one keyword.
+4.6 DELETE_ANNOTATION Learners delete their annotations, such as delete the highlight on one sentence.
+4.7 SEARCH_ANNOTATION Learners use search annotation tool to search and check their annotations.
+4.8 CLOSE_ANNOTATION Learners close the annotation window.
+5 CHATBOT 5.1 OPEN_CHATBOT Learners open the chatbot window.
+5.2 ASK_CHATBOT Learners type question and send questions to chatbot.
+5.3 READ_FEEDBACK_CHATBOT Learners move mouse, scroll mouse or select text in the chatbot response content.
+5.4 CLOSE_CHATBOT Learners close the chatbot window.
+6 NAVIGATION 6.1 PAGE_NAVIGATION Learners navigate through several pages (stay less than 6 seconds)
+6.2 TABLE_OF_CONTENT Learners check the table of content, such as scrolling in that area
+6.3 TRY_OUT_TOOLS Learners quickly (less than 3 seconds) open and close tools for the first time without using them
+7 OFF_TASK 7.1 OFF_TASK Learners being inactivity for more than 5 minutes
 Dialogue Annotation
@@
 Preference utterances solicit subjective opinions or recommendations, such as which idea might be easier or more appropriate.
-The remaining three subcategories—Provide Context, Revise, and Inspire—emerged inductively through close examination of our collected data.
+The remaining three subcategories—Provide Context, Revise, and Inspire—emerged inductively through close examination of the collected data.
 Provide Context captures utterances in which students offer background information or setup without making a direct information request, often to improve the relevance of future responses.
@@
 To ensure clarity and consistency in the annotation process, a detailed coding manual was developed, including operational definitions, decision rules, and illustrative examples for each code.
-This preparatory work was essential for aligning coders’ interpretations and reducing ambiguity during the annotation of both student intents and chatbot responses.
+Specifically, coders first labelled student intents and then labelled the corresponding chatbot responses.
+For student intents, precedence rules were applied: Inspire was used when the student explicitly extended the chatbot’s immediately preceding response (e.g., requesting more similar examples) without expressing dissatisfaction; Revise was used when the student indicated dissatisfaction with the prior response and reformulated the question, re-asked it with modifications, or added new constraints.
+When neither applied, coders distinguished Knowledge (requests for fixed, factual explanations/definitions) from Preference (requests for subjective opinions/recommendations, often framed as “Do you…?”) and Judgement (idea-seeking or evaluative requests with multiple valid answers, including imperatives such as “generate/give me…”).
+Writing-related requests were labelled as Task Clarification when asking about assignment requirements (e.g., what each section should include) and as Writing Advice when asking for editing, feedback, rewriting, or scoring of the student’s own text.
+Inputs that only provided isolated information or lacked a clear information need were labelled Provide Context, whereas greetings were labelled Chitchat and off-task prompts were labelled Others.
+For chatbot responses, coders labelled Fact for objective, well-defined answers (e.g., definitions), Suggestion for neutral guidance/options or generated content without evaluative judgement, and Critique when the response contained explicit evaluation or opinion; responses requesting additional user information were labelled Clarify, explicit inability to answer was labelled No Answer, and social/off-task replies were labelled Chitchat or Others accordingly.
+These rules complemented the codebook definitions and examples, ensuring consistent and reproducible labelling decisions across coders.
 The annotation process began with an inter-rater calibration phase, during which all three coders independently annotated a randomly selected subset of 400 utterances—comprising both student and chatbot turns—from the full dataset.
@@
 This dual-coding plus adjudication approach strengthened the reliability and validity of the final annotated dataset, supporting rigorous analysis of student–GenAI interactions in the context of IPS tasks.
-In total, the annotated dataset comprises 3,560 student prompts and chatbot responses, each labelled according to the developed codebooks described in Table 1 and Table 2, with the distribution of intent and response types visualised in Figure 1 and the distribution student activity visualised in Figure 2.
+In total, the annotated dataset comprises 3,560 interaction pairs (consisting of student prompts and corresponding chatbot responses), each labelled according to the developed codebooks described in Table 2 and Table 3, with the distribution of intent and response types visualised in Figure 1 and the distribution student activity visualised in Figure 2.
 Figure 1: Students Questions and Chatbot Answers Category Distribution
@@
 Data Records
-The dataset is publicly available on GitHub [^1] https://github.com/Xinyu-Li/FLoRA_Open_Data_Monash_FIT5145 under the BSD-2-Clause license.
-It comprises multiple CSV-format data files, as detailed below.
+The dataset is publicly available on GitHub [^1] https://github.com/Xinyu-Li/FLoRA_Open_Data_IPS under the BSD-2-Clause license. It comprises multiple CSV-format data files, as detailed below.
 Dialogue Data
@@
 Descriptions for all action categories can also be found in Table 1.
-Table 3: The learning action categories
-No. Action No. Sub_Action Sub Action Definition
-0 START 0.1 START_TASK Indicate when learners start the task.
-1 INSTRUCTION 1.2 TASK_REQUIREMENT Learners open task requirement page (task instruction page) to read about what the task requires them to do.
-2 READING 2.1 READING_MATERIALS Learners re-read relevant reading materials.
-2.2 REREADING_MATERIALS Learners re-read relevant reading materials.
-3 ESSAY 3.1 OPEN_ESSAY Learners open the essay and read their essay without new writing.
-3.2 WRITE_ESSAY Learners open the essay and write based on reading materials.
-3.3 PASTE_TEXT_ESSAY Learners copy and paste materials from reading content to the essay window.
-3.4 SAVE_ESSAY Learners click the save button to save the essay.
-3.5 CLOSE_ESSAY Learners close the essay window.
-4 ANNOTATION 4.1 CREATE_NOTE Learners create notes from reading materials.
-4.2 CREATE_HIGHLIGHT Learners create highlights on the reading materials.
-4.3 READ_ANNOTATION Learners click on annotations or open the annotation tool to read their notes or highlights.
-4.4 LABEL_ANNOTATION Learners label or add new labels, or accept suggested labels on their notes or highlights.
-4.5 EDIT_ANNOTATION Learners edit their annotations, such as edit one note on one keyword.
-4.6 DELETE_ANNOTATION Learners delete their annotations, such as delete the highlight on one sentence.
-4.7 SEARCH_ANNOTATION Learners use search annotation tool to search and check their annotations.
-4.8 CLOSE_ANNOTATION Learners close the annotation window.
-5 CHATBOT 5.1 OPEN_CHATBOT Learners open the chatbot window.
-5.2 ASK_CHATBOT Learners type question and send questions to chatbot.
-5.3 READ_FEEDBACK_CHATBOT Learners move mouse, scroll mouse or select text in the chatbot response content.
-5.4 CLOSE_CHATBOT Learners close the chatbot window.
-6 NAVIGATION 6.1 PAGE_NAVIGATION Learners navigate through several pages (stay less than 6 seconds)
-6.2 TABLE_OF_CONTENT Learners check the table of content, such as scrolling in that area
-6.3 TRY_OUT_TOOLS Learners quickly (less than 3 seconds) open and close tools for the first time without using them
-7 OFF_TASK 7.1 OFF_TASK Learners being inactivity for more than 5 minutes
-There are two types of web pages in FLoRA, namely instructions and inspirational_materials.
-The former refers to the only web page detailing the assignment instruction; the latter refers to various reading materials provided by the teachers for students to get inspiration about different topics.
 Writing logs
-The writing log data contains all the keystroke events and the writing content that happen while students writing the proposal. It contains the following columns:
+The writing logs enable writing-process analytics (e.g., drafting/revision dynamics, writing bursts, and temporal alignment between writing and reading/chatbot use), which is central to studying IPS synthesis/production with GenAI support.
+The writing log data contains all the keystroke events and the writing content that happen while students writing the proposal.
+It contains the following columns:
   • user_id - the unique identifier for each student.
@@
 Second, we tested the stability of the trace data across cohorts (temporal stability and replicability).
-We compared a pilot dataset collected under identical experimental conditions in Semester One 2024 (2024S1) with the main dataset from Semester Two 2024 (2024S2), which we intend to release as open data.
+As described in the Methods, the study was conducted in two offerings of the same course: a pilot cohort in 2024S1 and the main cohort in 2024S2 (released as open data).
+We assessed stability by comparing the distribution of the number of time-stamped raw trace events (log entries) per student between 2024S1 and 2024S2.
 After conducting normality and homogeneity‑of‑variance tests, we found that both datasets were non‑normal but exhibited equal variances.
@@
 This review ensured that anomalies or inappropriate items were identified and addressed promptly.
-Fourth, we validated the assessment data (assessment reliability).
-Expert teachers evaluated all submitted essays using detailed, criterion‑based rubrics.
-Their discipline expertise and consistent application of the criteria ensured rigorous grading and supported reliable links between learning outcomes and interactions with the GenAI system.
-Finally, we verified the survey data (background variable integrity).
-The research team reviewed all pre‑task and post‑task responses, removing incomplete entries when appropriate.
-This step ensured that the background information and participant feedback used in subsequent analyses were reliable.
-Together, these procedures triangulate platform validity, cross‑semester stability, dialogue quality, assessment reliability, and survey integrity, thereby strengthening the overall credibility of the dataset and the conclusions drawn from it.
+Fourth, we verified the assessment outcome measure (rubric transparency and integrity).
+As described in the Methods, students’ proposals were graded by experienced course teaching staff using a criterion-based rubric with predefined criteria and weights.
+Using an explicit rubric provides a structured and transparent learning-outcome measure for linking performance with students’ interactions and trace data in FLoRA.
+Finally, we verified survey data integrity.
+As described in the Methods, the research team reviewed pre-task and post-task survey responses and removed incomplete entries for key variables when appropriate.
+The complete survey instruments (all questions and response options) are provided in Supplementary Information Table S1, S2 and S3, enabling replication and reuse.
+Together, these procedures triangulate platform validity, cross-semester stability of trace data, dialogue content integrity, rubric-based outcome transparency, and survey data integrity, thereby strengthening the overall credibility of the dataset and the conclusions drawn from it.
 Usage Notes
-This dataset is best suited for research about IPS, as well as for detailed writing process analytics.
-Researchers can use the dialogues to infer planning, monitoring, evaluation, and adaptation processes across sessions.
-Because the interaction log data is pre‑labelled with SRL processes in FLoRA (using Bannert’s SRL framework [31, 29]), this data can also be used for SRL analysis to evaluate students cognition and metacognition activities during the learning.
+This dataset is best suited for research about IPS, including fine-grained analysis of students’ writing and revision processes during the synthesis/production stage while interacting with GenAI.
+Researchers can use the dialogue sequences—together with behavioural and writing logs—to examine indicators of planning, monitoring, evaluation, and adaptation across sessions.
+Because the interaction log data is pre-labelled with learning-action categories (Table 1), it can be used to analyse students’ IPS behaviours at a fine temporal granularity.
+Researchers interested in SRL process-level analysis can further map sequences of learning actions to Bannert-aligned cognitive and metacognitive SRL processes by applying an established process library (see Fan et al.
+[12]).
 Building on these process measures, the dataset also enables rich GenAI interaction studies and performance modelling.